Harbour Company makes two models of electronic tablets, the Home and the Work. Basic production information follows: Home WorRequired: 1. Suppose Harbour uses a traditional costing system with machine hours as the cost driver. Determine the amount of3. Calculate Harbours gross margin per unit for each product under the traditional costing system. (Round your intermediate5. Assuming an ABC system, assign overhead costs to each product based on activity demands. Overhead Assigned To Home Overhea6. Calculate the production cost per unit for each of Harbours products in an ABC system. (Round your intermediate calculati8. Compare the gross margin of each product under the traditional system and ABC. (Round your answers to 2 decimal places.) H
Answer #1
1. Overhead assigned to each product Home Work Total
a Total overhead 186180
b No. of Machine Hours 1300 1600 2900
c Cost per Machine hour (c=a/b) 64.20
d Overhead to be assigned (d= machine hour * cost per machine hour) 83460 102720 186180
e Expected production 620 470
f overhead cost per unit (f = d/e) 135 219
2. production cost per unit
Home Work
a Direct Materials 37 71
b Direct labor 15 37
c Overheads 134.61 218.55
d production cost per unit (sum from a to c) 186.61 326.55
3. Gross margin per unit
Home Work
a Sales price 363 578
b Production cost per unit 186.61 326.55
c production cost per unit (a - b) 176.39 251.45
4. activity Rate as per ABC
Activity cost Cost drivers Activity Rate per driver = (Activity cost /Cost drivers)
a Setup cost 86920 106 820
b Quality control 55760 680 82
c Maintenance 43500 2900 15
5. overhead cost to each product
Activity Rate per driver no. of drivers Total cost = a*b expected production cost per unit =c/d
Home a b c d e
Setup cost 820 38 31160 620 50.26
Quality control 82 340 27880 620 44.97
Maintenance 15 1300 19500 620 31.45
Work
Setup cost 820 68 55760 470 118.64
Quality control 82 340 27880 470 59.32
Maintenance 15 1600 24000 470 51.06
6. production cost per unit as per ABC
Home Work
a Direct Materials 37 71
b Direct labor 15 37
c Overheads
d Setup cost 50.26 118.64
e Quality control 44.97 59.32
f Maintenance 31.45 51.06
g production cost per unit (sum from a to f) 178.68 337.02
7. Gross margin per unit as per ABC
Home Work
a Sales price 363 578
b Production cost per unit 178.68 337.02
c production cost per unit (a - b) 184.32 240.98
8. comparison of Gross margin per unit
Home Work
a Traditional System 176.39 251.45
b ABC System 184.32 240.98
increase/decrease 7.94 -10.47